OpenStudy (ellamoyseyuk):

Evaluate the determinant for the following matrix:[1,4,4;5,2,2;1,5,5]

OpenStudy (ellamoyseyuk):

@agent0smith and @jim_thompson5910 can yall help me?

OpenStudy (agent0smith):

Screenshot.

OpenStudy (agent0smith):

http://www.purplemath.com/modules/determs2.htm

OpenStudy (ellamoyseyuk):

@agent0smith

OpenStudy (agent0smith):

idk i'm too tired to work it out, especially with you not posting it in an easy to read way.

jimthompson5910 (jim_thompson5910):

[1,4,4;5,2,2;1,5,5] would be written as \[\Large \begin{bmatrix}1 & 4 & 4\\5 & 2 & 2\\1 & 5 & 5\end{bmatrix}\]

OpenStudy (ellamoyseyuk):

i just did it the way that angent0smith told me to do and i got 0

jimthompson5910 (jim_thompson5910):

So we have this determinant (vertical bars signal "determinant") \[\Large \left|\begin{array}{ccc}1 & 4 & 4\\5 & 2 & 2\\1 & 5 & 5\end{array}\right|\]

OpenStudy (ellamoyseyuk):

okay

jimthompson5910 (jim_thompson5910):

Copy over the first two columns and write them on the right side of the determinant \[\Large \left|\begin{array}{ccc|cc}{\color{red}{1}} & {\color{green}{4}} & 4 & {\color{red}{1}} & {\color{green}{4}} \\{\color{red}{5}} & {\color{green}{2}} & 2 & {\color{red}{5}} & {\color{green}{2}}\\{\color{red}{1}} & {\color{green}{5}} & 5 & {\color{red}{1}} & {\color{green}{5}}\end{array}\right.\]

OpenStudy (ellamoyseyuk):

okay

jimthompson5910 (jim_thompson5910):

So we have this \[\Large \left|\begin{array}{ccc|cc}1 & 4 & 4 & 1 & 4\\5 & 2 & 2 & 5 & 2\\1 & 5 & 5 & 1 & 5\end{array}\right.\]

jimthompson5910 (jim_thompson5910):

Now what we do next is highlight the diagonals like shown in that link Let's go one diagonal at a time \[\Large \left|\begin{array}{ccc|cc}{\color{red}{1}} & 4 & 4 & 1 & 4\\5 & {\color{red}{2}} & 2 & 5 & 2\\1 & 5 & {\color{red}{5}} & 1 & 5\end{array}\right.\] What happens when you multiply those red values together?

OpenStudy (ellamoyseyuk):

=10

jimthompson5910 (jim_thompson5910):

correct, how about when you do this diagonal \[\Large \left|\begin{array}{ccc|cc}1 & {\color{red}{4}} & 4 & 1 & 4\\5 & 2 & {\color{red}{2}} & 5 & 2\\1 & 5 & 5 & {\color{red}{1}} & 5\end{array}\right.\]

OpenStudy (ellamoyseyuk):

=8

OpenStudy (ellamoyseyuk):

4*5*5=100

jimthompson5910 (jim_thompson5910):

\[\Large \left|\begin{array}{ccc|cc}1 & 4 & {\color{red}{4}} & 1 & 4\\5 & 2 & 2 & {\color{red}{5}} & 2\\1 & 5 & 5 & 1 & {\color{red}{5}}\end{array}\right.\] yes 4*5*5 = 100

OpenStudy (ellamoyseyuk):

1*2*4=8 5*2*1=10 5*5*4=100

jimthompson5910 (jim_thompson5910):

So far we have 10+8+100 = 118

OpenStudy (ellamoyseyuk):

10-8+8+100-10-100=0

jimthompson5910 (jim_thompson5910):

correct. The final result will be 0

OpenStudy (ellamoyseyuk):

okay thanks

jimthompson5910 (jim_thompson5910):

wolfram alpha confirms it http://www.wolframalpha.com/input/?i=determinant(%5B%5B1,4,4%5D,%5B5,2,2%5D,%5B1,5,5%5D%5D)